OpenStudy (anonymous):

am I correct or wrong on these 5 problems?

jigglypuff314 (jigglypuff314):

1, yep :) 2. 12x^3 + 5 + 5x^2 - 6x^2 + 3x^3 - 2 - x 5x^2 - 6x^2 = -x^2 not +x^2

OpenStudy (anonymous):

so would #2 be B?

jigglypuff314 (jigglypuff314):

2. B yep :) 3. yep :) 4. yep :) 5. foil (2h - 1)(3h+5) ^ ^ -1 * 5 = -5 not +5

OpenStudy (anonymous):

wow I really got 3 and 4 right? haha

OpenStudy (anonymous):

Would #5 be D?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Can you please help with 3 more? I'm not sure if I got them right or wrong @jigglypuff314

jigglypuff314 (jigglypuff314):

nope nope and nope lol gimme a sec for explanations

OpenStudy (anonymous):

aw I thought I got at least one of them right, but okay lol

jigglypuff314 (jigglypuff314):

21k^4 + 9k^3 - 12k^2 you want to pull out the greatest common factor so what do you think all three of these values have in common?

OpenStudy (anonymous):

3

jigglypuff314 (jigglypuff314):

yep and for the k values, you can only pull out the greatest common factor would of that too so what is the largest number of ks you can pull out?

OpenStudy (anonymous):

7

jigglypuff314 (jigglypuff314):

umm out of k^4 k^3 and k^2 you would only be able to pull out the one with the lowest power so k^2 then out of the overall value you can only pull out a 3k^2

OpenStudy (anonymous):

oops I read it wrong

OpenStudy (anonymous):

sorry

jigglypuff314 (jigglypuff314):

it's alright :)

jigglypuff314 (jigglypuff314):

the second is a difference of squares... a^2 - b^2 = (a+b)(a-b) 36x^2 - 25 = (6x)^2 - (5)^2 = then plug into form if 6x = a and 5 = b

OpenStudy (anonymous):

so would it be C?

jigglypuff314 (jigglypuff314):

yep :D

OpenStudy (anonymous):

yay! :D

jigglypuff314 (jigglypuff314):

x^2 - 5x + 6 ^ the numbers add up to -5 ^ the numbers multiply up to 6 for the option you choose (x + 2)(x - 3) ^ ^ 2 * -3 not = +6 nor do they add up 2 + -3 not = -5

OpenStudy (anonymous):

so would it be C?

jigglypuff314 (jigglypuff314):

it wouldn't be C because (x + 2)(x - 3) foiled out Does Not Equal x^2 - 5x + 6

OpenStudy (anonymous):

so it's either A or D, I'm not sure

jigglypuff314 (jigglypuff314):

x^2 - 5x + 6 ^ the numbers add up to -5 ^ the numbers multiply up to 6 I'm not really good at explaining how to factor polynomials :/

OpenStudy (anonymous):

It's okay, I'm not really good at understanding them

OpenStudy (anonymous):

I think I'll just guess and put D, idk

jigglypuff314 (jigglypuff314):

it's A I just don't know how to explain :P

OpenStudy (anonymous):

oh haha it's fine, what about the last one?

jigglypuff314 (jigglypuff314):

I'm confuzeled now... Use the polynomial to answer the question. 36x^2 - 25 Which represents the polynomial in factored form? A. (18x - 5)(18x - 5) B. (18x - 5)(18x + 5) C. (6x + 5)(6x - 5) <---- D. (6x - 5)(6x - 5) Use the polynomial to answer the question. x^2 - 5x + 6 Which represents the polynomial in factored form? A. (x - 2)(x - 3) <---- B. (x + 2)(x + 3) C. (x + 2)(x - 3) D. (x - 2)(x + 3)

OpenStudy (anonymous):

it was one of the 3 questions you said I got incorrect

jigglypuff314 (jigglypuff314):

foil the one you choose, see what you get.

OpenStudy (anonymous):

I'm not good at foil....

OpenStudy (anonymous):

I've been asking a lot of help today because I'm not in a quiet room and I haven't been in a quiet area all day because I currently do not have a room to stay in

jigglypuff314 (jigglypuff314):

practice makes perfect, and foil is a must (x + 2)(x - 3) = x^2 -3x + 2x - 6 = x^2 - x - 6

jigglypuff314 (jigglypuff314):

I don't mind helping ya :) it's a lot more fun working with you than with some other peoplez :P

OpenStudy (anonymous):

aw thank you :) you too, you don't get frustrated and you take your time with trying to help me understanding all the problems! its so nice

OpenStudy (anonymous):

I feel like the next best answer to put is D because it had - and +

jigglypuff314 (jigglypuff314):

you would want one that had the same signs cuz they have to multiply up to +6 if they had different signs you would end up with a -6 in the equation.

OpenStudy (anonymous):

So maybe it would be B...?

OpenStudy (anonymous):

They both have +'s

OpenStudy (anonymous):

I don't think it would be A because the signs are both -'s

jigglypuff314 (jigglypuff314):

but you would want a -5 in the middle part -2 -3 = -5 and -2*-3 = +6

jigglypuff314 (jigglypuff314):

- times - = +

OpenStudy (anonymous):

oh so it is A.

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Now I just have 3 more but if I'm honest, I didn't come up with an answer for any of them and I feel like I've been asking you a lot so I'll just out them in another question. UNLESS you'd want to try and see if they're easy for you. They are about factoring.

jigglypuff314 (jigglypuff314):

they'll be easy for me to do :P I'll be able to do them :) I just can't explain factoring, so yeah....

jigglypuff314 (jigglypuff314):

B C D

OpenStudy (anonymous):

Thank you so so so much! I wish I could give you more than a thank you and I've already gave you best response. I'm so greatful for your help and I now feel less stressed because of how kind and helpful you have been to me! You're amazing and intelligent! :D

jigglypuff314 (jigglypuff314):

lol your welcome I'm currently legitly failing my math class right now, so I'm not exactly smart, but I like simple math :) so feel free to tag or message me if you have any more questions :)

OpenStudy (anonymous):

I'm surprised to hear that as you're really good at math. You must be in a much higher math class than me because what you call simple math, I find really difficult lol. I will do but for today that's all I have to ask. :)

jigglypuff314 (jigglypuff314):

my math class lost me at "slope of a curve" -.-" anyways, glad I could help :) and good night ;D

OpenStudy (anonymous):

Good night! :D Hope you wake up in a good mood and have a good day tomorrow as you deserve it!